# Coherent control of optomechanical entanglement and steering via dual parametric amplification

### 16 Apr 2026 v1

@@ III Numerical simulations and discussions:
To further study the effect of thermal noise, we plot the entanglement and steering as functions of the environment temperature $T$ in the presence of coherent feedback. Fig. 7 shows that, within the stability region, properly tuning the reflectivity can effectively improve the thermal robustness. However, as can be seen from Fig. 7(a), when the system is tuned excessively close to the stability threshold, the enhancement of reflectivity has a destructive effect on the entanglement.
IV Experimental Feasibility And Applications:
@@
The OPA has been well developed in both theoretical studies and experimental implementations [51, 56, 57, 58, 59, 60, 61, 62]. It can be realized by placing a PPKTP crystal inside the cavity and driving it with a classical pump field [63, 64, 65]. And the effective strength can be tuned by the pump power, whereas the corresponding squeezing phase can be adjusted through the relative phase locking among the pump, signal, and local-oscillator fields [63, 64].
@@
To detect and verify the entanglement and steering, it is inevitable to measure the elements of reduced covariance matrix [68, 69]. The cavity field quadratures can be directly accessed by homodyne or heterodyne detection of the cavity output [70, 71]. The mechanical quadratures cannot be directly measured but can be inferred with the help of an auxiliary probe cavity mode which is weak enough to be negligible. When the probe field, with frequency $\omega_{p,l}$, is red detuned from the resonance frequency $\omega_{p,c}$ of the auxiliary probe cavity by the mechanical frequency $\omega_{b}$, i.e., $\omega_{p,l}\approx\omega_{p,c}-\omega_{b}$, the linearized interaction between the probe cavity mode and the mechanical mode reduces to the beam-splitter form, which maps the mechanical state onto the probe output field. In this way, the CM elements associated with the mechanical mode can be reconstructed from the correlations of the probe output field [69, 72, 73].
Besides, the feedback delay $t_{d}$ is an important factor that should be addressed. As we mentioned in Appendix. B, the instantaneous feedback approximation is valid when $|2\kappa_{a}r_{b}t_{d}|\ll 1$. For a feedback loop of length about $10~\mathrm{cm}$, the propagation delay is only of order $10^{-10}$$10^{-9}~\mathrm{s}$, which is much shorter than the cavity field lifetime $1/\kappa_{a}$. Therefore, the feedback delay can be safely neglected, and the coherent feedback can be treated as instantaneous. Also, if the delay is not negligible, imprinting a phase on light in the feedback loop is a feasible approach to cancel the effect of delay [44].
V Conclusion:
In conclusion, we have investigated the role of coherent feedback in regulating the quantum correlations and the local mechanical quantum properties. Due to introducing coherent feedback, the effective decay rate and the corresponding matrix elements of the photon mode in the diffusion matrix are modulated. The coherent feedback in our system can enhance the intracavity photon and make the entanglement stronger and its parameter region broader with the joint modulation of OPA and MPA. Besides, the coherent feedback enlarges the parameter region supporting one way steering and enables the transition from one-way to two-way steering. Moreover, the system can sustain relatively high purity while exhibiting strong squeezing, thereby improving the quality of the local nonclassical resource state. Besides, we also find that compared to the system only modulated by OPA and MPA, our entanglement and steering is more robust against the thermal noises. Our work may provide a more controllable quantum resource, which have potential applications in continuous-variable quantum communication and quantum key distribution.
VI Acknowledgments:
@@ Appendix A The limit-cycle dynamics
$\dot{\alpha}=-\bigl(\kappa_{\mathrm{fb}}+i\Delta_{\mathrm{fb}}\bigr)\alpha+ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+tE,$ (16)
@@
$\frac{d}{dt}|\alpha|^{2}=\alpha^{\ast}\dot{\alpha}+\alpha\dot{\alpha}^{\ast}=-2\kappa_{\mathrm{fb}}(r)|\alpha|^{2}+2\,\mathrm{Re}\!\left[\alpha^{\ast}\Bigl(ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+t_{b}E\Bigr)\right].$ (17)
In the long-time regime the system settles into a time-periodic limit cycle. Averaging Eq. (17) over one period $T$ of the limit cycle, we obtain the power-balance relation
$2\kappa_{\mathrm{fb}}(r)\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}=P_{\mathrm{om}}(r)+P_{\mathrm{OPA}}(r)+P_{\mathrm{drv}},$ (18)
where the right-hand side corresponds to the cycle-averaged power injected into the cavity by the optomechanical interaction, by the OPA and by the external drive, respectively.
For the parameter set used in Fig. 2 we have $g,G_{c}\ll\kappa_{\mathrm{fb}}$, so that the drive term $P_{\mathrm{drv}}$ dominates, while $P_{\mathrm{om}}$ and $P_{\mathrm{OPA}}$ provide only small corrections and depend only weakly on $r$. In this experimentally relevant limit Eq. (18) yields the approximate scaling
$\bigl\langle|\alpha|\bigr\rangle_{T}\propto\frac{1}{\kappa_{\mathrm{fb}}(r)}=\frac{1}{\kappa(1-2r\cos\theta)},\qquad$ (19)
@@ Appendix B Validity of the instantaneous-feedback approximation
To justify the instantaneous-feedback approximation used in the main text, we explicitly restore the finite propagation delay in the coherent-feedback loop. To avoid confusion with the Floquet period, we denote the feedback delay time by $t_{d}$ and the loop phase by $\phi_{f}$. The feedback-modified input field is then written as
$\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}(t)=r_{b}e^{\mathrm{i}\phi_{f}}a_{\mathrm{out}}(t-t_{d})+t_{b}a_{\mathrm{in}}(t).$ (20)
@@
$a_{\mathrm{out}}(t-t_{d})=\sqrt{2\kappa_{a}}\,a(t-t_{d})-a_{\mathrm{in}}(t-t_{d}),$ (21)
@@
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (22)
@@
$\Delta_{a}^{\prime}=\Delta_{a}-g\bigl(\langle b\rangle+\langle b\rangle^{\ast}\bigr).$ (23)
@@
$\delta a(t-t_{d})=\delta a(t)-t_{d}\,\delta\dot{a}(t)+\mathcal{O}(t_{d}^{2}).$ (24)
Substituting Eq. (24) into Eq. (22) yields
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (25)
$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\left[\delta a(t)-t_{d}\,\delta\dot{a}(t)\right]+\mathrm{noise}.$
@@
$\left[1+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}t_{d}\right]\delta\dot{a}(t)$ $=\left[-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\right]\delta a(t)$ (26)
$\quad+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+\mathrm{noise}.$
@@
$\left|2\kappa_{a}r_{b}t_{d}\right|\ll 1.$ (27)

### 18 Apr 2026 v2

@@ 3 Numerical simulations and discussions
To further study the effect of thermal noise, we plot the entanglement and steering as functions of the environment temperature $T$ in the presence of coherent feedback. Fig. 7 shows that, within the stability region, properly tuning the reflectivity can effectively improve the thermal robustness. However, as can be seen from Fig. 7(a), when the system is tuned excessively close to the stability threshold, the enhancement of reflectivity has a destructive effect on the entanglement. Compared with the scheme in Ref. [15], the quantum correlations are more robust against thermal noise.
4 Experimental Feasibility And Applications
@@
The OPA has been well developed in both theoretical studies and experimental implementations [54, 60, 61, 62, 63, 64, 65, 66]. It can be realized by placing a PPKTP crystal inside the cavity and driving it with a classical pump field [67, 68, 69]. The effective strength can be tuned by the pump power, whereas the corresponding squeezing phase can be adjusted through the relative phase locking among the pump, signal, and local-oscillator fields [67, 68].
@@
To detect and verify the entanglement and steering, it is necessary to measure the elements of the reduced CM[72, 73]. The cavity field quadratures can be directly accessed by homodyne or heterodyne detection of the cavity output [74, 75]. The mechanical quadratures cannot be directly measured but can be inferred with the help of an auxiliary probe cavity mode which is weak enough to be negligible. When the probe field, with frequency $\omega_{p,l}$, is red detuned from the resonance frequency $\omega_{p,c}$ of the auxiliary probe cavity by the mechanical frequency $\omega_{b}$, i.e., $\omega_{p,l}\approx\omega_{p,c}-\omega_{b}$, the linearized interaction between the probe cavity mode and the mechanical mode reduces to the beam-splitter form, which maps the mechanical state onto the probe output field. In this way, the CM elements associated with the mechanical mode can be reconstructed from the correlations of the probe output field [73, 76, 77].
Besides, the feedback delay $t_{d}$ is an important factor that should be addressed. As we mentioned in Appendix. B, the instantaneous feedback approximation is valid when $|2\kappa_{a}r_{b}t_{d}|\ll 1$. For a feedback loop of length about $10~\mathrm{cm}$, the propagation delay is only of order $10^{-10}$$10^{-9}~\mathrm{s}$, which is much shorter than the cavity field lifetime $1/\kappa_{a}$. Therefore, the feedback delay can be safely neglected, and the coherent feedback can be treated as instantaneous. Also, if the delay is not negligible, imprinting a phase on light in the feedback loop is an feasible approach to cancel the effect of delay [45].
5 Conclusion
In conclusion, we investigated the synergistic effect of the coherent feedback and parametric amplifications in regulating the quantum correlations and the local mechanical quantum properties. Our investigation shows that the coherent feedback has two roles: the modulation of the effective decay rate and the corresponding matrix elements in the diffusion matrix of the photon mode. Besides, the coherent feedback enlarges the parameter region supporting one-way steering and enables the transition from one-way to two-way steering. The frequency and amplitude ratio play an important role in enhancing and controlling the entanglement. Moreover, the system can sustain relatively high purity while exhibiting strong squeezing, thereby improving the quality of the local nonclassical resource state. Modulating the phase shift provides an effective way in controlling quantum correlations. Finally, we find that the quantum correlations in the COM system are more robust against thermal noise compared to the system only modulated by the OPA and MPA. Our work may provide a more controllable route to engineering quantum resource, with potential applications in continuous-variable quantum communication and quantum key distribution.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).
Acknowledgments
@@ Appendix A The limit-cycle dynamics
$\dot{\alpha}=-\bigl(\kappa_{\mathrm{fb}}+i\Delta_{\mathrm{fb}}\bigr)\alpha+ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+tE,$ (A1)
@@
$\frac{d}{dt}|\alpha|^{2}=\alpha^{\ast}\dot{\alpha}+\alpha\dot{\alpha}^{\ast}=-2\kappa_{\mathrm{fb}}(r)|\alpha|^{2}+2\,\mathrm{Re}\!\left[\alpha^{\ast}\Bigl(ig\,\alpha(\beta+\beta^{\ast})+2G_{c}\alpha^{\ast}e^{-i(\Omega_{c}t-\theta_{c})}+t_{b}E\Bigr)\right].$ (A2)
In the long-time regime the system settles into a time-periodic limit cycle. Averaging Eq. (A2) over one period $T$ of the limit cycle, we obtain the power-balance relation
$2\kappa_{\mathrm{fb}}(r)\,\bigl\langle|\alpha|^{2}\bigr\rangle_{T}=P_{\mathrm{om}}(r)+P_{\mathrm{OPA}}(r)+P_{\mathrm{drv}},$ (A3)
where the right-hand side corresponds to the cycle-averaged power injected into the cavity by the optomechanical interaction, by the OPA, and by the external drive, respectively.
For the parameter set used in Fig. 2 we have $g,G_{c}\ll\kappa_{\mathrm{fb}}$, so that the drive term $P_{\mathrm{drv}}$ dominates, while $P_{\mathrm{om}}$ and $P_{\mathrm{OPA}}$ provide only small corrections and depend only weakly on $r$. In this experimentally relevant limit Eq. (A3) yields the approximate scaling
$\bigl\langle|\alpha|\bigr\rangle_{T}\propto\frac{1}{\kappa_{\mathrm{fb}}(r)}=\frac{1}{\kappa(1-2r\cos\theta)},$ (A4)
@@ Appendix B Validity of the instantaneous-feedback approximation
To justify the instantaneous-feedback approximation used in the main text, we explicitly restore the finite propagation delay in the coherent-feedback loop. To avoid confusion with the Floquet period, we denote the feedback delay time by $t_{d}$ and the loop phase by $\phi_{f}$. The feedback modified input field is then written as
$\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}(t)=r_{b}e^{\mathrm{i}\phi_{f}}a_{\mathrm{out}}(t-t_{d})+t_{b}a_{\mathrm{in}}(t),$ (B1)
@@
$a_{\mathrm{out}}(t-t_{d})=\sqrt{2\kappa_{a}}\,a(t-t_{d})-a_{\mathrm{in}}(t-t_{d}),$ (B2)
@@
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (B3)
@@
$\Delta_{a}^{\prime}=\Delta_{a}-g\bigl(\langle b\rangle+\langle b\rangle^{\ast}\bigr).$ (B4)
@@
$\delta a(t-t_{d})=\delta a(t)-t_{d}\,\delta\dot{a}(t)+\mathcal{O}(t_{d}^{2}).$ (B5)
Substituting Eq. (B5) into Eq. (B3) yields
$\delta\dot{a}(t)$ $=-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)\delta a(t)+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]$ (B6)
$\quad+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\left[\delta a(t)-t_{d}\,\delta\dot{a}(t)\right]+\mathrm{noise},$
@@
$\left[1+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}t_{d}\right]\delta\dot{a}(t)$ $=\left[-\left(\mathrm{i}\Delta_{a}^{\prime}+\kappa_{a}\right)+2\kappa_{a}r_{b}e^{\mathrm{i}\phi_{f}}\right]\delta a(t)$ (B7)
$\quad+\mathrm{i}G_{\mathrm{eff}}\bigl[\delta b(t)+\delta b^{\dagger}(t)\bigr]+2G_{c}e^{-\mathrm{i}(\Delta_{c}t-\theta_{c})}\delta a^{\dagger}(t)+\mathrm{noise},$
@@
$\left|2\kappa_{a}r_{b}t_{d}\right|\ll 1.$ (B8)